Physics 335—Thermal Physics Homework Assignment #6 Due Friday, October 18, 2024, 11:40xam Note the unusual due date.

Ch 5:

Problems 5.6, 5.10, 5.13, and 5.20 (20 pts.xeach).

Hints: For Problem 5.13, you will need one of the results from Problem 5.12, which we will discuss in class.

Extra Credit:

Physics Colloquium: Wednesday, October 16, at 4:30pm in Hugel 100. Dr. Fei Huang from the Weizmann Institute of Science will talk about “Dark-Matter Archaeology: Reconstructing the History of Our Universe from the Present-Day Distribution of Matter.” (20 pts.)